# Designing quantum chemistry algorithms with just-in-time compilation

### v5

@@ 3.2 Quartet fragmentation (1qnt) algorithm
For high-angular momentum integrals, the intermediate variables and integral elements of one shell quartet cannot fit in the registers. From GPU4PySCF v1.0-v1.4, a fixed-size 1D fragmentation strategy is introduced to split the integral into multiple threads. Instead of one block thread processing all integrals of one shell quartet, a group of 2 to 256 threads cooperatively compute all integral elements within one shell quartet. The strategy significantly improved the performance of the high-angular integrals. In this work, we introduce the dynamic fragmentation strategy to further improve the data locality, and multi-level reduction algorithm to reduce the bottleneck of memory access (both shared memory and global memory). Since multiple threads for the same shell quartet use the same intermediate variables, the intermediate variable $I_{x}$, $I_{y}$ and $I_{z}$ will be stored in shared memory. Then, each thread in the group will compute and store into registers only a fraction of the integrals. For example, the shape of the integral (dd|dd) is (6,6,6,6), each thread computes a fragment with the size of (3,3,3,3). 16 threads are needed to store the entire (dd|dd) integral. In the reduction stage, each thread is responsible for a fragment of 3x3 $v_{j}$ and $v_{k}$. The fragments $v_{j}$ and $v_{k}$ are further reduced at the thread block level. The reduction scheme is presented in Algorithm 1. The shape of the intermediate variables are also known at compilation time. A similar idea has been proposed in [3], where $(dd|dd)$ is partitioned in $i$, $j$, and $k$ indices. $(ff|ff)$ is partitioned in $i$ and $j$ indices.
@@
2: //Distribute fragments to different threads
3:for $f_{i}=0$ to $N^{f}_{i}/F_{i}$ in parallel do
4:  for $f_{j}=0$ to $N^{f}_{j}/F_{j}$ in parallel do
5:   for $f_{k}=0$ to $N^{f}_{k}/F_{k}$ in parallel do
6:    for $f_{l}=0$ to $N^{f}_{l}/F_{l}$ in parallel do
7:     // Integral fragment $T_{ijkl}=(ij|kl)$
8:     $T^{\text{reg}}[\colon N^{f}_{i},\colon N^{f}_{j},\colon N^{f}_{k},\colon N^{f}_{l}]=0$
9:     for $i_{p}=0$ to $n_{i}^{p}$ do
10:      for $j_{p}=0$ to $n_{j}^{p}$ do
11:       for $k_{p}=0$ to $n_{k}^{p}$ do
12:        for $l_{p}=0$ to $n_{l}^{p}$ do
13:         // Compute integrals with Eq. (1)-(3)
14:         $T^{\text{reg}}[i,j,k,l]\mathrel{+}=c_{p}c_{q}c_{r}c_{s}(pq|rs)$
15:        end for
@@
19:     Read $D_{J}^{\text{reg}}[:F_{k},:F_{l}]$ and $D_{K}^{\text{reg}}[:F_{j},:F_{l}]$
20:     $J^{\text{reg}}[\colon F_{i},\colon F_{j}]=0,K^{\text{reg}}[\colon F_{i},\colon F_{k}]=0$
21:     for $i=0$ to $F_{i}$ do
22:      for $j=0$ to $F_{j}$ do
23:       for $k=0$ to $F_{k}$ do
24:        for $l=0$ to $F_{l}$ do
25:         //Thread level reduction
26:         $J^{reg}[i,j]\mathrel{+}=\sum_{kl}T^{\text{reg}}[i,j,k,l]D_{J}^{\text{reg}}[k,l]$
27:         $K^{reg}[i,k]\mathrel{+}=\sum_{jl}T^{\text{reg}}[i,j,k,l]D_{K}^{\text{reg}}[j,l]$
28:        end for
@@
32:     //Block level reduction
33:     $J^{\text{shd}}[iF_{i}\!:\!(i+1)F_{i},\,jF_{j}\!:\!(j+1)F_{j}]\mathrel{+}=J^{\text{reg}}[\!:\!F_{i},\!:\!F_{j}]$
34:     $K^{\text{shd}}[iF_{i}\!:\!(i+1)F_{i},\,kF_{k}\!:\!(k+1)F_{k}]\mathrel{+}=K^{\text{reg}}[\!:\!F_{i},\!:\!F_{k}]$
35:    end for
@@
39: atomicAdd($J[:\!N^{f}_{i},:\!N^{f}_{j}],J^{\text{shd}}[:\!N^{f}_{i},:\!N^{f}_{j}]$)
40: atomicAdd($K[:\!N^{f}_{i},:\!N^{f}_{k}],K^{\text{shd}}[:\!N^{f}_{i},:\!N^{f}_{k}]$)
Optimize the fragmentation scheme. The performance of the integral kernel is sensitive to the fragment size. Different fragmentation strategies can easily lead to a 10x difference in performance. The fragment size can impact the register usage of each thread, the shared memory usage of each thread block, and the number of shell quartet assigned to each block. The resource consumption also need to meet the hard requirements of CUDA compute capability. We have to take these factors into account for searching the optimal fragment size. This optimization problem is less likely to be solved analytically. We look for the optimal fragmentation strategy by grid search. The searching space of high angular momentum can be huge. We made several assumptions:
@@
  • The optimal fragment size is independent of the number of primitives. This assumption in general is not true, since the kernel with more number of primitives take more registers. Then, less registers are left for the integral fragment. The assumption is made for compromising with the high computational cost. We search the optimal fragment size based on a fixed number of primitives for low-angular orbital (s, p orbitals). The number of primitives for high-angular atomic orbital (d, f, g orbitals) is set to be 1.
@@
Since the double-precision and single-precision algorithms differ in their usage of shared memory, register count, and hardware units, a strategy optimized for double precision is not necessarily optimal for single precision. Therefore, we optimize fragmentation strategies separately for each precision. The default fragmentation schemes are tuned on NVIDIA A100-80G and A10-24G GPUs, respectively. To ensure compatibility, the fragment size is optimized under the constraint of 48 KB shared memory per streaming multiprocessor, allowing the strategy to be applicable to older GPU generations. An optimization script is included in the JoltQC release, enabling users to retune the fragment size for different GPU architectures if they diverge significantly from the default targets.
We benchmark the 1tnq algorithm with a straight peptide chain composed of 30 glycines (gly30) and def2-TZVPP [31] basis set. Density matrix screening is turned off. Results on NVIDIA A100-80G and A10-24G GPUs appear in Figure 4 and 5, respectively. Using the default fragmentation for FP64, the 1tnq algorithm is 5-8x faster than GPU4PySCF v1.4 on A100-80G. When switched to FP32, the algorithm is further accelerated by 2x. For high-angular-momentum integrals, the computational cost of exp and erf functions is not significant relative to multiply/add operations. On A10-24G, the high-angular-momentum integrals are accelerated by roughly 2x for FP64. This is due to the algorithm in GPU4PySCF v1.4 being memory bound. The 1qnt algorithm reduces the memory bottleneck by multi-level reduction. When switched to FP32, the integrals are accelerated by 10-20x. Further performance analysis is provided in Appendix (Section 8).
@@ 4.1 Compilation Overhead
Just-in-time compilation introduces non-negligible overhead, but this cost is amortized through on-disk caching and in-memory reuse. When a kernel is first invoked, it is compiled and the resulting binary is saved to disk. The cached binaries can be reused across the process. During the first SCF iteration, each process loads these binaries into RAM directly; all subsequent iterations then fetch them directly from memory. Table 3 summarizes the one-time compilation times for various kernels using CUDA v12.1. As expected, high–angular-momentum integrals take longer to compile and generate larger binaries: for the C, H, O, and N elements with the 6-31G* basis (140 kernels), the compilation time is just 30 s and the total kernel size is roughly 10 MB, while the def2-TZVPP basis (750 kernels) yields about 100 MB of binary and takes approximately 200 s to compile. Once those kernels are compiled, subsequent calls take less than 1 s to load cached binaries.
These figures serve as a reference. Actual compilation times will vary with different CUDA versions and hardware quality, particularly disk performance. On modern devices, the compilation overhead can be reduced in various ways. In practice, especially in homogeneous cloud environments, it can be advantageous to store compiled kernels on shared or remote storage and download them as needed. The similar approach has been used in the machine learning community, e.g. PyTorch AOTInductor. Alternatively, retaining intermediate .ptx files and loading them directly allows users to bypass one step of compilation.
Integral pattern Algorithm
Compilation
time (ms)
Binary size
@@ 4.2 Benchmark with GPU4PySCF and TeraChem
In this work, integrals for the Coulomb (J) and exchange (K) terms are screened simultaneously. GPU4PySCF and JoltQC use the same screening technique. The number of significant shell quartets in each JK calculation is identical in both cases. In double-precision, JoltQC reproduces GPU4PySCF’s results exactly, so we omit a detailed FP64 accuracy analysis. Under our single precision scheme, JK matrix elements are computed in single precision but accumulated in FP64; all other operations including diagonalization and direct inversion of the iterative subspace (DIIS)—remain in FP64.
We use a customized TeraChem v1.9 build with f-orbital support. Although both GPU4PySCF and TeraChem employ the Schwarz inequality to bound integral magnitudes, they apply it at different levels: TeraChem at the primitive shell level and GPU4PySCF at the contracted level. Their SCF initial guesses and DIIS settings also differ, so our GPU4PySCF–TeraChem comparisons are not strictly apples-to-apples. For all calculations shown, the screening threshold for both J and K matrices is set to $10^{-13}$. SCF with GPU4PySCF/JoltQC converges in 14 iterations, whereas the same SCF with TeraChem converges in 11 iterations. This does not imply TeraChem’s settings are superior, as convergence behavior can vary with the system. A sample TeraChem input is provided in Appendix 7. All the basis sets are in Cartesian format.
When both codes run in FP64, their total energies agree to within 0.001 mHa. When all integrals are evaluated in FP32, we observe uncertainties on the order of 1 mHa in both JoltQC and TeraChem (Table 4), reinforcing the well-known result that pure single-precision arithmetic is insufficient for high-accuracy quantum-chemical calculations. A more practical mixed-precision algorithm is among our future directions.
Molecule Basis
GPU4PySCF/FP64
- JoltQC/FP32
GPU4PySCF/FP64
- TeraChem/FP64
JoltQC/FP32
- TeraChem/FP32
@@
Table 4: The accuracy of Hartree-Fock (HF) energy with JoltQC v0.1 against TeraChem v1.9. GPU4PySCF/FP64 and JoltQC/FP64 matches exactly and is not shown in the table.
As shown in Table 5, JoltQC delivers a twofold speedup over GPU4PySCF v1.4 when using the 6-31G* basis set and a fourfold gain with the larger def2-TZVPP basis. Although TeraChem v1.9 generally outpaces GPU4PySCF v1.4, its advantage is only marginal. In double-precision, JoltQC outperforms TeraChem by about 4x when def2-TZVPP basis set is used, and a less significant 1.6x for 6-31G* basis. In single-precision, JoltQC slightly outperforms TeraChem on the small basis set and runs twice as fast on def2-TZVPP. Both TeraChem and JoltQC still employ double precision for eigenvalue decomposition and DIIS, making these linear-algebra routines a significant cost on the A10-24 GB GPU. Isolating the JK kernel, JoltQC is approximately three times faster than TeraChem.
@@
gly30 HF/6-31G*/A100/FP64 1.9 s 28.1 s 1.3 s 19.1 s 0.8 s 10.4 s
gly30 HF/def2-TZVPP/A100/FP64 78.4 s 1,125.7 s 91.4 s 1,055.1 s 18.8 s 268.1 s
gly30 HF/6-31G*/A10/FP32 - - 1.0 s 23.1 s 0.7 s 16.5 s
gly30 HF/def2-TZVPP/A10/FP32 - - 63.5 s 870.8 s 20.0 s 406.1 s
Table 5: Benchmark among GPU4PySCF v1.4, TeraChem v1.9, and JoltQC v0.1, on A100-80G using FP64 arithmetic, and on A10-24G using FP32 arithmetic.

### v6

@@ 3.2 Quartet fragmentation (1qnt) algorithm
For high-angular-momentum integrals, the intermediate variables and integral elements of one shell quartet cannot fit in the registers. From GPU4PySCF v1.0-v1.4, a fixed-size 1D fragmentation strategy was introduced to split the integral into multiple threads. Instead of a single thread processing all integrals of one shell quartet, a group of 2 to 256 threads cooperatively compute all integral elements within one shell quartet. The strategy significantly improved the performance of the high-angular-momentum integrals. In this work, we introduce the dynamic fragmentation strategy to further improve the data locality, and multi-level reduction algorithm to reduce the bottleneck of memory access (both shared memory and global memory). Since multiple threads for the same shell quartet use the same intermediate variables, the intermediate variables $I_{x}$, $I_{y}$ and $I_{z}$ are stored in shared memory. Then, each thread in the group will compute and store into registers only a fraction of the integrals. For example, for $(dd|dd)$, the integral tensor has shape (6,6,6,6). With fragment size (3,3,3,3), each dimension is split into 6/3 = 2 parts. The four fragment indices ($f_{i}$, $f_{j}$, $f_{k}$, $f_{l}$) are distributed across threads, requiring $2^{4}=16$ threads per shell quartet. After computing its integral fragment $T^{reg}$, each thread contracts $T^{reg}$ with the density matrix to produce local contributions to J and K. We denote the thread-local J contribution as $v_{j}$ (shape $F_{i}\times F_{j}$) and the thread-local K contribution as $v_{k}$ (shape $F_{i}\times F_{k}$). These are accumulated in registers during the thread-level reduction (Algorithm 1, lines 21–30), then reduced across threads at the block level (lines 32–36).
The reduction scheme is presented in Algorithm 1. The shape of the intermediate variables are also known at compilation time. A similar idea has been proposed in [3], where $(dd|dd)$ is partitioned in $i$, $j$, and $k$ indices, while $(ff|ff)$ is partitioned in $i$ and $j$ indices.
@@
1: // Superscripts: reg = per-thread registers, shd = per-block shared memory, no superscript = global memory.
@@
3: //Distribute fragments to different threads
4:for $f_{i}=0$ to $N^{f}_{i}/F_{i}$ in parallel do
5:  for $f_{j}=0$ to $N^{f}_{j}/F_{j}$ in parallel do
6:   for $f_{k}=0$ to $N^{f}_{k}/F_{k}$ in parallel do
7:    for $f_{l}=0$ to $N^{f}_{l}/F_{l}$ in parallel do
8:     // Integral fragment $T_{ijkl}=(ij|kl)$
9:     $T^{\text{reg}}[\colon\!N^{f}_{i},\colon\!N^{f}_{j},\colon\!N^{f}_{k},\colon\!N^{f}_{l}]=0$ // Allocate integral fragment in registers
10:     for $i_{p}=0$ to $n_{i}^{p}$ do
11:      for $j_{p}=0$ to $n_{j}^{p}$ do
12:       for $k_{p}=0$ to $n_{k}^{p}$ do
13:        for $l_{p}=0$ to $n_{l}^{p}$ do
14:         // Compute integrals with Eq. (1)-(3)
15:         $T^{\text{reg}}[i,j,k,l]\mathrel{+}=c_{ip}c_{jq}c_{kr}c_{ls}(pq|rs)$
@@
19:     end for
20:     Read $D_{J}^{\text{reg}}[:\!F_{k},:\!F_{l}]$ and $D_{K}^{\text{reg}}[:\!F_{j},:\!F_{l}]$ // Read density matrix subblocks in registers
21:     $J^{\text{reg}}[\colon\!F_{i},\colon\!F_{j}]=0,K^{\text{reg}}[\colon\!F_{i},\colon\!F_{k}]=0$ // Allocate accumulators for J, K in registers
22:     for $i=0$ to $F_{i}$ do
23:      for $j=0$ to $F_{j}$ do
24:       for $k=0$ to $F_{k}$ do
25:        for $l=0$ to $F_{l}$ do
26:         //Thread level reduction
27:         $J^{reg}[i,j]\mathrel{+}=\sum_{kl}T^{\text{reg}}[i,j,k,l]D_{J}^{\text{reg}}[k,l]$
28:         $K^{reg}[i,k]\mathrel{+}=\sum_{jl}T^{\text{reg}}[i,j,k,l]D_{K}^{\text{reg}}[j,l]$
@@
32:     end for
33:     //Block level reduction
34:     $J^{\text{shd}}[iF_{i}\!:\!(i+1)F_{i},\,jF_{j}\!:\!(j+1)F_{j}]\mathrel{+}=J^{\text{reg}}[\!:\!F_{i},\!:\!F_{j}]$
35:     $K^{\text{shd}}[iF_{i}\!:\!(i+1)F_{i},\,kF_{k}\!:\!(k+1)F_{k}]\mathrel{+}=K^{\text{reg}}[\!:\!F_{i},\!:\!F_{k}]$
@@
39:end for
40: atomicAdd($J[:\!N^{f}_{i},:\!N^{f}_{j}],J^{\text{shd}}[:\!N^{f}_{i},:\!N^{f}_{j}]$) // Write block results to global J
41: atomicAdd($K[:\!N^{f}_{i},:\!N^{f}_{k}],K^{\text{shd}}[:\!N^{f}_{i},:\!N^{f}_{k}]$) // Write block results to global K
Optimize the fragmentation scheme. The performance of the integral kernel is sensitive to the fragment size. Different fragmentation strategies can easily lead to a 10x difference in performance. The fragment size can impact the register usage of each thread, the shared memory usage of each thread block, and the number of shell quartet assigned to each block. The resource consumption also need to meet the hard requirements of CUDA compute capability. We have to take these factors into account for searching the optimal fragment size. This optimization problem is unlikely to be solved analytically. We look for the optimal fragmentation strategy by grid search. For each candidate fragmentation, we compile and execute the kernel on a representative workload and select the configuration with the lowest measured runtime. The searching space of high angular momentum can be huge. We made several assumptions:
@@
  • The optimal fragment size is independent of the number of primitives. This assumption in general is not true, since the kernel with more number of primitives take more registers. Then, less registers are left for the integral fragment. This assumption is highly empirical. In actual runs, kernels with more primitives consume more registers, leaving fewer for the integral fragment. However, jointly optimizing over both fragment size and primitive count would make the search space intractable. We therefore fix the primitive count at a representative value and search only over the space of fragment size. For high-angular-momentum shells ($d$, $f$, $g$), which typically have fewer primitives, we set the primitive count to 1.
@@
We optimize fragmentation strategies separately for the combinations of precisions and GPU devices. As previously mentioned, the FP64 algorithm may use a different fragmentation scheme from the FP32 algorithm for the same angular momentum. Since the double-precision and single-precision algorithms differ in their usage of shared memory, register count, and hardware units, a strategy optimized for double precision is not necessarily optimal for single precision. The default fragmentation schemes are tuned on NVIDIA A100-80G and A10-24G GPUs, respectively. To ensure compatibility, the fragment size is optimized under the constraint of 48 KB shared memory per streaming multiprocessor, allowing the strategy to be applicable to older GPU generations. An optimization script is included in the JoltQC release, enabling users to retune the fragment size for different GPU architectures if they diverge significantly from the default targets.
We benchmark the 1qnt algorithm with a straight peptide chain composed of 30 glycines (gly30) and def2-TZVPP [32] basis set. Density matrix screening is turned off. Results on NVIDIA A100-80G and A10-24G GPUs appear in Figure 4 and 5, respectively. Using the default fragmentation for FP64, the 1qnt algorithm is 5-8x faster than GPU4PySCF v1.4 on A100-80G. When switched to FP32, the algorithm is further accelerated by 2x. For high-angular-momentum integrals, the computational cost of exp and erf functions is not significant relative to multiply/add operations.
On A10-24G, the high-angular-momentum integrals are accelerated by roughly 2x for FP64. This is due to the algorithm in GPU4PySCF v1.4 being memory bound. The 1qnt algorithm reduces the memory bottleneck by multi-level reduction. When switched to FP32, the integrals are accelerated by 10-20x. Further performance analysis is provided in Appendix (Section 8).
@@ 4 Results
We evaluate JoltQC along three axes: compilation overhead (Sec. 4.1), end-to-end SCF performance against GPU4PySCF v1.4 and TeraChem v1.9 (Sec. 4.2), and single-precision accuracy. All benchmarks use a single NVIDIA A100-80G GPU for FP64 and an A10-24G GPU for FP32 unless otherwise noted.
@@ 4.1 Compilation Overhead
Just-in-time compilation introduces non-negligible overhead, but this cost is amortized through on-disk caching and in-memory reuse. When a kernel is first invoked, it is compiled and the resulting binary is saved to disk. The cached binaries can be reused across the processes. During the first SCF iteration, each process loads these binaries into RAM directly; all subsequent iterations then fetch them directly from memory. Table 3 summarizes the one-time compilation times for various kernels using CUDA v12.1. As expected, high–angular-momentum integrals take longer to compile and generate larger binaries: for the C, H, O, and N elements with the 6-31G* basis (140 kernels), the compilation time is just 30 s and the total kernel size is roughly 10 MB, while the def2-TZVPP basis (750 kernels) yields about 100 MB of binary and takes approximately 200 s to compile. Once those kernels are compiled, subsequent calls take less than 1 s to load cached binaries.
These figures serve as a reference. Actual compilation times will vary with different CUDA versions and hardware quality, particularly disk performance. In practice, the compilation overhead can be reduced in various ways. Especially in homogeneous cloud environments, it can be advantageous to store compiled kernels on shared or remote storage and download them as needed. A similar approach is used in the machine learning community, where PyTorch’s AOTInductor pre-compiles GPU kernels and caches the resulting binaries for reuse across runs. Alternatively, retaining the intermediate PTX assembly (a CUDA specific compilation intermediate) and loading it directly allows users to skip the final machine-code generation step on subsequent invocations.
Integral pattern Algorithm Compilation time (ms) Binary size (KB)
@@ 4.2 Benchmark with GPU4PySCF and TeraChem
In this work, integrals for the Coulomb (J) and exchange (K) terms are screened simultaneously. GPU4PySCF and JoltQC use the same screening technique. The number of significant shell quartets in each JK calculation is identical in both cases. In double precision, JoltQC reproduces GPU4PySCF’s J and K matrices to machine precision for all test systems, confirming that the JIT-compiled kernels are numerically identical to the AOT implementation. We therefore focus the accuracy analysis on the single-precision results. Under our single precision scheme, JK matrix elements are computed in single precision but accumulated in FP64; all other operations including diagonalization and direct inversion of the iterative subspace (DIIS)—remain in FP64.
We use a customized TeraChem v1.9 build with f-orbital support. Before presenting timings, we note several differences that affect the fairness of comparisons. GPU4PySCF and JoltQC use identical screening (Schwarz inequality at the contracted level), so their JK kernel timings are directly comparable. TeraChem applies Schwarz screening at the primitive shell level, uses different SCF initial guesses and DIIS settings, and converges in fewer iterations (e.g. 11 vs. 14 for the gly30 molecule). Consequently, GPU4PySCF/JoltQC–TeraChem comparisons reflect differences in both algorithms and settings, and are not strictly apples-to-apples. For all calculations, the screening threshold is set to $10^{-13}$. SCF with GPU4PySCF/JoltQC converges in 14 iterations, whereas the same SCF with TeraChem converges in 11 iterations. This does not imply TeraChem’s settings are superior, as convergence behavior can vary with the system. A sample TeraChem input is provided in Appendix 7. All the basis sets are in Cartesian format.
When both codes run in FP64, their total energies agree to within 0.001 mHa. When all integrals are evaluated in FP32, we observe uncertainties on the order of 1 mHa in both JoltQC and TeraChem (Table 4), reinforcing the well-known result that pure single-precision arithmetic is insufficient for high-accuracy quantum-chemical calculations. For the cc-pVQZ basis, which involves higher angular momentum, the FP32 energy deviations remain sub-milliHartree (0.16 mHa for tamoxifen and 0.23 mHa for sphingomyelin), indicating that single-precision accuracy does not degrade with increasing basis set size. A more practical mixed-precision algorithm is among our future directions.
Molecule Basis GPU4PySCF/FP64 - JoltQC/FP32 GPU4PySCF/FP64 - TeraChem/FP64 JoltQC/FP32 - TeraChem/FP32
@@
Tamoxifen $c$-pVQZ 0.16 mHa - -
Sphingomyelin $c$-pVQZ 0.23 mHa - -
Table 4: The accuracy of Hartree-Fock (HF) energy with JoltQC v0.1 against TeraChem v1.9. GPU4PySCF/FP64 and JoltQC/FP64 matches exactly and is not shown in the table. Dashes indicate that the software does not support the required angular momentum (g functions in $c$-pVQZ). Each column reports the total energy difference $E_{A}-E_{B}$ between the two codes listed.
As shown in Table 5, JoltQC delivers a twofold speedup over GPU4PySCF v1.4 when using the 6-31G* basis set and a fourfold gain with the larger def2-TZVPP basis. Although TeraChem v1.9 generally outpaces GPU4PySCF v1.4, its advantage is only marginal. In double-precision, JoltQC outperforms TeraChem by about 4x when the def2-TZVPP basis set is used, and a less significant 1.6x for the 6-31G* basis. The advantage grows further with larger basis sets. For the cc-pVQZ basis, which includes g-type functions not supported by TeraChem v1.9, JoltQC achieves a 7–8× speedup over GPU4PySCF v1.4 on the tamoxifen, raffinose, and sphingomyelin systems — nearly doubling the gain observed with def2-TZVPP. This trend confirms that JIT compilation becomes increasingly beneficial as the angular momentum of the basis set rises, precisely the regime where register pressure and loop complexity penalize statically compiled kernels the most. In single-precision, JoltQC slightly outperforms TeraChem on the small basis set and runs twice as fast on def2-TZVPP. Both TeraChem and JoltQC still employ double precision for eigenvalue decomposition and DIIS. On the A10-24G, which has only 1/32 the FP64 throughput of its FP32 units, these FP64 linear-algebra routines consume a disproportionate fraction of the total SCF time, diminishing the effective speedup from FP32 integral evaluation. Isolating the JK kernel on the A10-24G with FP32 arithmetic, JoltQC is approximately three times faster than TeraChem across the benchmark systems (Table 5, JK/cycle column).
@@
gly30 6-31G*/A100/FP64 1.9 s 28.1 s 1.3 s 19.1 s 0.8 s 10.4 s
gly30 def2-TZVPP/A100/FP64 78.4 s 1,125.7 s 91.4 s 1,055.1 s 18.8 s 268.1 s
gly30 6-31G*/A10/FP32 - - 1.0 s 23.1 s 0.7 s 16.5 s
gly30 def2-TZVPP/A10/FP32 - - 63.5 s 870.8 s 20.0 s 406.1 s
valinomycin 6-31G*/A100/FP64 2.4 s 34.5 s 1.9 s 24.2 s 1.1 s 14.8 s
valinomycin def2-TZVPP/A100/FP64 113.0 s 1,579.6 s 162.0 s 1,641.6 s 30.0 s 419.2 s
valinomycin $c$-pVTZ/A100/FP64 96.0 s 1,349.5 s 140.0 s 1,534.5 s 24.0 s 330.6 s
tamoxifen $c$-pVQZ/A100/FP64 152.0 s 1,977.4 s - - 21.4 s 278.9 s
raffinose $c$-pVQZ/A100/FP64 188.0 s 2,252.9 s - - 28.6 s 343.6 s
sphingomyelin $c$-pVQZ/A100/FP64 227.0 s 2,726.0 s - - 29.7 s 356.1 s
Table 5: Benchmark among GPU4PySCF v1.4, TeraChem v1.9, and JoltQC v0.1, on A100-80G using FP64 arithmetic, and on A10-24G using FP32 arithmetic. Dashes indicate that the software does not support the required angular momentum (g functions in $c$-pVQZ)